OpenStudy (anonymous):

Factor and find all solutions of 3x^3 - 6x^2 - 9x = 0

OpenStudy (anonymous):

\[3x^3 - 6x^2 - 9x = 0\]

jimthompson5910 (jim_thompson5910):

you can start off by factoring out the GCF 3x 3x^3 - 6x^2 - 9x = 0 3x(x^2 - 2x - 3) = 0 now factor x^2 - 2x - 3

OpenStudy (anonymous):

So that would be 3x(x-1)^2 = 3

jimthompson5910 (jim_thompson5910):

no you can't move that 3 over like that

jimthompson5910 (jim_thompson5910):

and x^2 -2x doesn't factor to (x-1)^2

OpenStudy (anonymous):

I used the completing the square method...

jimthompson5910 (jim_thompson5910):

factor x^2 - 2x - 3 find two numbers that multiply to -3 and add to -2

OpenStudy (anonymous):

I don't think there r any...

OpenStudy (anonymous):

This is how i did it, \[3x^3-6x-9x=0\] \[3x(x^2-2x-2)=0\] \[3x(x^2-2x+1)= 2+1\] \[3x(x^2-2x+1)=3\] \[3x(x-1)^2 =3\] \[3x(x-1)=\pm \sqrt{3}\] \[3x(x)=1\pm \sqrt{3}\]

OpenStudy (anonymous):

Then i got stuck there

jimthompson5910 (jim_thompson5910):

sure there are: -3 and +1

jimthompson5910 (jim_thompson5910):

-3 + 1 = -2 -3 * 1 = -3

jimthompson5910 (jim_thompson5910):

since the two numbers are -3 and 1, this means x^2 - 2x - 3 factors to (x-3)(x+1)

jimthompson5910 (jim_thompson5910):

3x^3 - 6x^2 - 9x = 0 3x(x^2 - 2x - 3) = 0 3x(x-3)(x+1) = 0 now use the zero product property

OpenStudy (anonymous):

but how did u get the 3 because u get a radical if you find the square root of 3

jimthompson5910 (jim_thompson5910):

in your steps, you factored 3x^3 - 6x^2 - 9x incorrectly

jimthompson5910 (jim_thompson5910):

you should get 3x(x^2 - 2x - 3) NOT 3x(x^2 - 2x - 2)

jimthompson5910 (jim_thompson5910):

also, I'm not completing the square. I'm simply factoring x^2 - 2x - 3

OpenStudy (anonymous):

ooooohhhhhhh

OpenStudy (anonymous):

Thanks lol it was just a silly mistake

jimthompson5910 (jim_thompson5910):

you're welcome